OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (–2, 1)?

OpenStudy (anonymous):

M=(6-1)/(3--2)=5/5--->M=1 Y-6=1(x-3)---->Y-6=x-3--->ANSWER IS Y=X+3

OpenStudy (anonymous):

I meant to say in point slope that would be y-6=1(x-3)